OpenStudy (anonymous):

Find the slope for the table of values below. x 1 2 3 4 y 8 12 16 20 @jojo4eva

OpenStudy (jojo4eva):

so it wants the slope for the whole table of data

OpenStudy (anonymous):

yup

OpenStudy (lucaz):

slope is dy/dx (change in y over change in x), for every unit of x you increase 4 units of y, so the slope is m=4/1=4

OpenStudy (anonymous):

so what would the answer be with these choices y = 5x + 5 y = 7x + 5 y = 5x + 2 y = 7x + 22

OpenStudy (anonymous):

no sry wrong answers

OpenStudy (anonymous):

4 8 1/8 1/4

OpenStudy (jojo4eva):

@lucaz can u help her please thanks cuz im kinda clueless on this one

OpenStudy (lucaz):

x=1, y=4, m=dy/dx=4/1=4

OpenStudy (jojo4eva):

ok i got it now lol i was just confused on which data i was finding the slope of

OpenStudy (anonymous):

O! ok i i get it now! THANK YOU!!

OpenStudy (lucaz):

no problem